OpenStudy (anonymous):

Reopening question; Hey there, I need help with this distance btw 2 points.

OpenStudy (anonymous):

\[(\sqrt{3}, \sqrt{2}) and (\frac{ 2 }{ \sqrt{3}} ,-\sqrt{2}) \]

OpenStudy (anonymous):

Some helped me and got this answer but I got confused about his answer and was too late to ask him \[\sqrt{(-\sqrt2-\sqrt2)^2+(\frac2{\sqrt3}-\sqrt3)^2} =\sqrt{8+\frac13} =\sqrt{\frac{75}9}=\frac{5\sqrt3}3\]

OpenStudy (anonymous):

in this what is yr cofusiin

OpenStudy (anonymous):

How did he get the second step.

OpenStudy (anonymous):

let me telll

OpenStudy (anonymous):